OpenStudy (cutiecomittee123):

A forensic detective is called to the scene of a murder at 2am when she checks the temperature of the body, she finds that it is 80 degrees Fahrenheit. The temperature of the room in which the body is found is 65 degrees Fahrenheit. If the detective uses Newtons law of cooling T(t)=Ta+(To-Ta)e^-0.1947t , what will she say was the time of death? So far I wrote t(0)=65-144e^-0.1947(0) Next I need help with these two questions An archeologist finds an artifact that contains 44% of the carbon-14 that it would have had origionally.

OpenStudy (cutiecomittee123):

(the half life of carbon-14 is 5700 years)

OpenStudy (wolf1728):

I'll take question 2 I'm guessing you need to know how much time has passed. There's formulas and a calculator here: http://www.1728.org/halflife.htm but I'll give you the answer. elapsed time = half-life * [log (begng amt / endg amt)]/log(2) we'll say begng amt is 100% and endng amt is 44% elapsed time = 5,700 * [log (100 / 44)]/log(2) elapsed time = 5,700 * [log 2.2727272727] / log(2) elapsed time = 5,700 * [0.35654732351 / 0.30102999566] elapsed time = 5,700 * 1.1844245711 elapsed time = 6,751.22 years

OpenStudy (cutiecomittee123):

Awesome thank you :))